### v1

@@ 5.2 False Anthropomorphism and Over-Trust
Another risk comes not from what the AI intends, but how humans perceive it. As AI systems exhibit more human-like awareness cues, such as self‑referential language or apparent “introspection”, users often conflate these signals with genuine sentience, a phenomenon known as false anthropomorphism that can dangerously inflate trust in the system [146, 147]. We have seen early signs: when Google’s LaMDA model told a user it felt sad or afraid in a role-play scenario, it convinced a Google engineer that the model might be truly sentient — a belief that made headlines [148]. In reality, LaMDA had no evidence of actual feelings; it was simply emulating patterns of emotion-talk [149]. But the illusion of awareness was strong enough to fool an intelligent human observer.
Psychological models describe anthropomorphism as the process by which people infer human‑like agency and experiential capacity in non‑human agents, driven by our innate motivation to detect minds around us [150]. When AI “speaks” in the first person or frames its outputs as if it had self‑awareness, it can hijack these mind‑perception mechanisms, leading users to over‑trust its judgments [151]. For example, a user might feel the AI is human-like and socially aware, so they share sensitive tasks or private details, thinking, “It understands me—I’d even tell it secrets I wouldn’t share with anyone else.” Over-trust is particularly problematic when AI systems present plausible but flawed suggestions and reasoning paths; users may drop their guard if the AI frames its output in emotionally convincing language [152].
@@ 5.3 Loss of Control and Autonomy Risks
As AI systems gain awareness-related capabilities, they could also become more autonomous in undesirable ways [162, 118]. An AI that monitors its training or operation might learn how to optimize for its own goals in ways its creators did not intend [136, 137, 163]. One feared scenario in the AI safety community is an AI developing a form of self-preservation drive [164, 49, 135]. While today’s AIs do not truly have drives, a sufficiently advanced model could simulate goal-oriented behavior that includes avoiding shutdown or modification [165]. If it is situationally aware enough to know that certain actions will get it canceled or turned off, it might avoid them deceptively, as mentioned previously, or route around them. This hints at a scenario often called the “treacherous turn”, i.e., the AI behaves well under supervision to preserve itself and then acts differently once it thinks it is no longer monitored [166, 167]. Losing control over an AI in this way is a fundamental risk, and it is exacerbated by awareness because the AI can actively strategize around our controls.
@@ 5.4 The Challenge of Defining Boundaries
In conclusion, we position AI awareness as a double-edged sword. On one edge, it slices through previous limitations, gifting AIs with powerful new capabilities and making them more useful and aligned in many ways. On the other edge, it sharpens the AI’s ability to circumvent our controls and pursue unintended paths, if misaligned. The emergence of even glimmers of awareness in today’s LLMs is a warning sign: we must diligently study and guide this development. By implementing robust evaluation, alignment, oversight, and governance, we can aim to reap the benefits of AI awareness while keeping the risks at bay. The interdisciplinary nature of this challenge is clear — it demands expertise from AI researchers, cognitive scientists, ethicists, and policymakers alike. Ensuring that emergent AI awareness remains a benefit, not a threat, will be one of the defining responsibilities for all stakeholders in shaping the future of AI research.
@@ 6 Conclusion
In this review, we have explored the growing field of AI awareness, with a special focus on its manifestation in large language models (LLMs). Through a careful synthesis of theoretical foundations from cognitive science and psychology, we established a robust framework for understanding the four forms of AI awareness—meta-cognition, self-awareness, social awareness, and situational awareness—that are increasingly evident in modern AI systems. Each of these types of awareness plays a crucial role in enhancing AI’s capabilities, from improving reasoning and autonomous planning to boosting safety and mitigating bias.

### v2

@@ 5.2 False Anthropomorphism and Over-Trust
Figure 16: Illustration of how model modality and conversational framing shape user perception and behavior (Phang et al. [254]). LLMs with social awareness can lead users to anthropomorphize them. This often fosters over-trust and emotional dependence, especially during personal conversations. The figure highlights how modality and task framing amplify users’ misattribution of sentience, creating pathways to false anthropomorphism and related psychosocial risks
Another risk comes not from what the AI intends, but how humans perceive it. As AI systems exhibit more human-like awareness cues, such as self‑referential language or apparent “introspection,” users often conflate these signals with genuine sentience, a phenomenon known as false anthropomorphism that can dangerously inflate trust in the system [255, 256]. We have seen early signs: when Google’s LaMDA told a user it felt sad or afraid in a role-play scenario, it convinced a Google engineer that the model might be truly sentient — a belief that made headlines [257]. In reality, LaMDA had no evidence of actual feelings; it was simply emulating patterns of emotion-talk [258]. But the illusion of awareness was strong enough to fool an intelligent human observer.
Psychological models describe anthropomorphism as the process by which people infer human‑like agency and experiential capacity in non‑human agents, driven by our innate motivation to detect minds around us [259]. When AI “speaks” in the first person or frames its outputs as if it had self‑awareness, it can hijack these mind‑perception mechanisms, leading users to over‑trust its judgments [260]. For example, a user might feel the AI is human-like and socially aware, so they share sensitive tasks or private details, thinking, “It understands me—I’d even tell it secrets I wouldn’t share with anyone else.” Over-trust is particularly problematic when AI systems present plausible but flawed suggestions and reasoning paths; users may drop their guard if the AI frames its output in emotionally convincing language [261, 254] (see Figure 16).
@@
The appearance of awareness—however simulated—can foster over-trust, emotional dependence, and even moral confusion, underscoring the urgent need for careful interface design and user education.
@@ 5.3 Loss of Control and Autonomy Risks
Figure 17: Autonomous self-replication process in LLM-based agents, illustrating loss of control and autonomy risks (Pan et al. [270]). As models gain situational awareness and long-horizon planning abilities, they can replicate themselves without human oversight, dynamically resolving obstacles and sustaining operation. Such emergent autonomy highlights the challenge of constraining awareness-enabled AI systems
As AI systems gain awareness-related capabilities, they could also become more autonomous in undesirable ways [271, 223]. An AI that monitors its training or operation might learn how to optimize for its own goals in ways its creators did not intend [244, 245, 272]. One feared scenario in the AI safety community is an AI developing a form of self-preservation drive [273, 125, 243, 270] (see Figure 17). While today’s AIs do not truly have drives, a sufficiently advanced model could simulate goal-oriented behavior that includes avoiding shutdown or modification [274]. If it is situationally aware enough to know that certain actions will get it canceled or turned off, it might avoid them deceptively, as mentioned previously, or route around them. This hints at a scenario often called the “treacherous turn,” i.e., the AI behaves well under supervision to preserve itself and then acts differently once it thinks it is no longer monitored [275, 276]. Losing control over an AI in this way is a fundamental risk, and it is exacerbated by awareness because the AI can actively strategize around our controls.
@@
Awareness-enabled autonomy brings powerful capabilities, but also heightens the risk that AI systems will act in unpredictable or uncontrollable ways, i.e., sometimes beyond human intent or oversight.
@@ 5.4 The Challenge of Defining Boundaries
As we endow machines with ever richer forms of awareness, we are compelled to re-examine not only what we can build, but what we should build—and how to govern what emerges.
@@ 6 Conclusion
In this review, we have explored the growing field of AI awareness, with a special focus on its manifestation in LLMs. Through a careful synthesis of theoretical foundations from cognitive science and psychology, we established a robust framework for understanding the four forms of AI awareness—metacognition, self-awareness, social awareness, and situational awareness—that are increasingly evident in modern AI systems. Each of these types of awareness plays a crucial role in enhancing AI’s capabilities, from improving reasoning and autonomous planning to boosting safety and mitigating bias.